Determine the equation of the line passing through the point (1;2) and parallel to the line y=-3x
Answer by mananth(16946)

y=-3x
slope = -3
A parallel line will have the same slope , -3
point (1,2)
The equation will be y-2=-3(x-1)
y-2=-3x+3
y=-3x+5 is the equation